OpenStudy (anonymous):

Assume that all variables are positive. Part A: Show that . Part B: Simplify . Part C: If , what is the value of h in terms of g? (going to post equations soon)

OpenStudy (anonymous):

\[part A:\sqrt{\sqrt{x}}=\sqrt[4]{x}\]\[part B:\sqrt[4]{\sqrt[4]{x}}\]\[part C: \sqrt[g]{\sqrt[g]{x}}=\sqrt[h]{x}\]

OpenStudy (misty1212):

i wonder what "show" means here...

OpenStudy (misty1212):

if you use exponents you get it, but that is just a trick of notation

OpenStudy (misty1212):

\[\sqrt{\sqrt{x}}=\left(x^{\frac{1}{2}}\right)^{\frac{1}{2}}\] \[=x^{\frac{1}{2}\times \frac{1}{2}}\] \[=x^{\frac{1}{4}}\] \[=\sqrt[4]{x}\]

OpenStudy (misty1212):

same for B \[\frac{1}{4}\times\frac{1}{4}=\frac{1}{16}\] giving you \[\sqrt[16]{x}\]

OpenStudy (danjs):

\[\sqrt{x}=x ^{1/2}~~~~~so~~~~\sqrt{\sqrt{x}}=[x ^{1/2}]^{1/2} = x ^{1/2 * 1/2} = x ^{1/4}\]

OpenStudy (danjs):

lol you beat me to it.

OpenStudy (misty1212):

and for the last one, just like the first two \[g^2=h\]

OpenStudy (danjs):

I think the point of this is to realize the property \[\sqrt[n]{x} = x ^{1/n}\]

OpenStudy (danjs):

then the normal 3 or 4 exponent properties

OpenStudy (danjs):

\[\sqrt[g]{\sqrt[g]{x}} = [\sqrt[g]{x}]^{1/g} = [x ^{1/g}]^{1/g} = x ^{1/g^2}~~~~~and\] \[\sqrt[h]{x} = x ^{1/h}\] setting equal \[x ^{1/g^2} = x ^{1/h}~~~~so~~~~\frac{ 1 }{ h^2 } = \frac{ 1 }{ g }\]

OpenStudy (danjs):

sorry the h and g are mistyped in the very last expression... it is \[\frac{ 1 }{ g^2 }=\frac{ 1 }{ h }~~~~so~~~~g^2 = h\]